OpenStudy (anonymous):

find the vector v the same direction with u=(-2,4) and magnitude of v=6

OpenStudy (anonymous):

Hint: describe the vector v as (x, y): then you can find x and y from the following equations: A) x^2 + y^2=36 B) y/x=-4/2=-2